OpenStudy (anonymous):

Solve the proportion. 5/7= x/35

OpenStudy (anonymous):

Multiply both sides with 35 35(5/7)=35( x/35) 25=x x=25

OpenStudy (anonymous):

25?

OpenStudy (anonymous):

Thank youu .

OpenStudy (anonymous):

25

OpenStudy (anonymous):

welcome :D